# Finite-time convergence to an $\epsilon$-efficient Nash equilibrium in potential games

### v3

@@ Appendix A Markov chains and mixing times
[…]
where in line $4$ we used that $\log(1+x)\leq x$ for $x>-1$ and then in line $6$ we used that for any $\mu_{0}$, $D(\mu^{0}:\mu)\leq\log\frac{1}{\mu_{\min}}$ for $\mu_{\min}:=\min_{a\in\mathcal{A}^{N}}\mu(a)$. Thus, we conclude that the mixing time is upper-bounded as follows:
$t_{\operatorname{mix}}^{P}(\epsilon)\leq\frac{1}{\rho(PP^{*})}\bigg{(}\log\log\frac{1}{\mu_{\min}}+2\log\frac{1}{\epsilon}\bigg{)}.$ (15)
Appendix B Convergence of log-linear learning
We first prove Lemma 3.3 and then state a novel lemma which we will use to prove Lemma 3.4. Lastly, we provide a formal proof of the convergence time of log-linear learning in Theorem 3.1.
B.1 Proof of Theorem 3.1
We are now ready to proceed with the proof of Theorem 3.1.
Proof B.1.
of Theorem 3.1 By Lemma 3.4, the log-Sobolev constant $\rho(PP^{*})$ can be lower-bounded as:
$\rho(PP^{*})\geq\frac{16\pi^{2}A^{N}\mu_{\min}p_{\min}^{3}}{25N^{2}A^{2}}\geq\frac{16\pi^{2}e^{-4\beta}}{25N^{2}A^{5}},$ (16)
where we used that by definition of $P$ in (4) and $\mu$ in (5) $\mu_{\min}$ and $p_{\min}$ can be lower-bounded as follows:
$\mu_{\min}=\min_{a\in\mathcal{A}^{N}}\mu(a)\geq\frac{e^{-\beta}}{A^{N}}$
Equation (2) in Lemma 2.2 provides an upper bound on the mixing time:
Plugging the bound on the log-Sobolev constant from Equation (16) into the equation above we obtain:
$t_{\operatorname{mix}}^{P}(\epsilon/2)$ $\leq\frac{25N^{2}A^{5}}{16\pi^{2}}e^{4\beta}\bigg{(}\log\log\frac{1}{\mu_{\min}}+2\log\frac{4}{\epsilon}\bigg{)}$
$\leq\frac{25N^{2}A^{5}}{16\pi^{2}}e^{4\beta}\bigg{(}\log\log\frac{A^{N}}{e^{-\beta}}+2\log\frac{4}{\epsilon}\bigg{)}$
$\leq\frac{25N^{2}A^{5}}{16\pi^{2}}e^{4\beta}\bigg{(}\log\log A^{N}+\log\beta+2\log\frac{4}{\epsilon}\bigg{)}.$
By setting $t\geq\frac{25N^{2}A^{5}}{16\pi^{2}}e^{4\beta}\bigg{(}\log\log A^{N}+\log\beta+2\log\frac{4}{\epsilon}\bigg{)}$ and $\beta$ as in Equation (6) we obtain:
$\mathbb{E}_{a\sim\mu^{t}}[\Phi(a)]$ $\geq\mathbb{E}_{a\sim\mu}[\Phi(a)]-2\|\mu^{t}-\mu\|_{TV}\max_{a\in\mathcal{A}^{N}}\Phi(a)$
$=\max_{a\in\mathcal{A}^{N}}\Phi(a)-\epsilon,$
where the third line follows from Lemma 3.3, the fact that $\|\mu^{t}-\mu\|_{TV}\leq\epsilon/2$, and the fact that $\Phi(\cdot)\in[0,1]$. This concludes the proof of Theorem 3.1.
B.2 Proof of Lemma 3.3
Proof B.2.
[…]
$=\sum_{a\in\mathcal{A}^{N}_{*}}\frac{\Phi(a)}{\sum_{\tilde{a}\in\mathcal{A}^{N}(\epsilon)}e^{\beta(\Phi(\tilde{a})-\Phi(a))}+\sum_{\tilde{a}\in\mathcal{A}^{N}\setminus\mathcal{A}^{N}(\epsilon)}e^{\beta(\Phi(\tilde{a})-\Phi(a))}}$
$\geq\sum_{a\in\mathcal{A}^{N}_{*}}\frac{\Phi(a)}{\sum_{\tilde{a}\in\mathcal{A}^{N}(\epsilon)}e^{0}+\sum_{\tilde{a}\in\mathcal{A}^{N}\setminus\mathcal{A}^{N}(\epsilon)}e^{-\beta\min\{\epsilon,\Delta\}}}$
$\geq\frac{A^{N}_{*}}{A^{N}(\epsilon)+(A^{N}-A^{N}(\epsilon))e^{-\beta\min\{\epsilon,\Delta\}}}\Phi(a^{*}),$ (17)
where the suboptimality gap $\Delta$ is given by $\Delta:=\min_{a\in\mathcal{A}^{N}:\Phi(a)<\Phi(a^{*})}\left(\Phi(a^{*})-\Phi(a)\right)$ with $a^{*}\in\mathcal{A}^{N}_{*}$. Then, we have that:
$\beta\geq\frac{1}{\max\{\epsilon,\Delta\}}$ $\log\bigg{(}(A^{N}-A^{N}(\epsilon))\bigg{(}\frac{1}{\epsilon A^{N}(\epsilon)}-\frac{1}{A^{N}(\epsilon)}\bigg{)}\bigg{)}$
[…]
By injecting the last inequality into Equation (17) we deduce that for $\beta\geq\frac{1}{\max\{\epsilon,\Delta\}}\log(A^{N}-A^{N}(\epsilon))(\frac{1}{\epsilon A^{N}(\epsilon)}-\frac{1}{A^{N}(\epsilon)}))$, it holds that:
$\mathbb{E}_{a\sim\mu}[\Phi(a)]\geq(1-\epsilon)\max_{a\in A^{N}}\Phi(a)\geq\max_{a\in A^{N}}\Phi(a)-\epsilon,$
where we used that $\Phi(a)\leq 1$ for all $a\in\mathcal{A}^{N}$.
B.3 Proof of Lemma 3.4
We first state the following lemma which will be used to prove Lemma 3.4.
Lemma B.3.
Consider the Markov chain $\{\hat{X}_{t}\}_{t\in\mathbb{N}}$ over a finite state space $\mathcal{A}^{N}$ with transition matrix $\hat{P}\in\mathbb{R}^{A^{N}\times A^{N}}$ given by:
$\hat{P}_{a,\tilde{a}}=\frac{1}{NA}\mathds{1}_{\tilde{a}\in\mathcal{N}(a)}$ (18)
where $\mathcal{N}(a)=\{\tilde{a}\in\mathcal{A}^{N}\>|\>\exists i\in[N]:\tilde{a}_{-i}=a_{-i}\}$. Assuming $A\geq 4$, the log-Soblev constant of $\hat{P}$ is lower-bounded by:
$\rho(\hat{P})\geq\frac{4\pi^{2}}{25NA}.$
Proof B.4.
We will lower bound the log-Soblev constant of $\hat{P}$ in terms of the log-Soblev constant of another Markov chain for which a lower bound on the log-Soblev constant is known.
First, note that the Markov chain $\{\hat{X}_{t}\}_{t\in\mathbb{N}}$ is aperiodic and irreducible with stationary distribution $\hat{\mu}(a)=1/A^{N}$. This can be verified by checking the detailed balance equations given by $\hat{\mu}(a)\hat{P}_{a,\tilde{a}}=\hat{\mu}(\tilde{a})\hat{P}_{\tilde{a},a}$ for all $a,\tilde{a}\in A^{N}$. Next, we make use of Corollary $2.15$ in (Montenegro et al., 2006) to lower-bound $\rho(\hat{P})$ in terms of the log-Soblev constant of another Markov chain $\bar{X}_{t}$ with transition matrix $\bar{P}$ and stationary distribution $\bar{\mu}$. Namely,
$\rho(\hat{P})\geq\frac{1}{MC}\rho(\bar{P}),$
where
$M=\max_{a\in\mathcal{A}^{N}}\frac{\hat{\mu}(a)}{\bar{\mu}(a)},$
$C=\max_{a\neq\tilde{a}:\hat{P}_{a,\tilde{a}}\neq 0}\frac{\bar{\mu}(a)\bar{P}_{a,\tilde{a}}}{\hat{\mu}(a)\hat{P}_{a,\tilde{a}}}.$
As the comparison Markov chain, we consider the product chain $\{\bar{X}_{t}\}_{t\in\mathbb{N}}$ with $\bar{X}_{t}=\prod_{i=1}^{N}\bar{X}_{i,t}$ on the state space $\mathbb{Z}_{K^{N}}=\prod_{i=1}^{N}\mathbb{Z}_{K}$. Here, each $\{\bar{X}_{i_{t}}\}_{t\in\mathbb{N}}$ is a simple random walk on the finite circle $\mathbb{Z}_{K}=\{1,\ldots,K\}$ with $K\geq 4$. The transition matrix and the stationary distribution of the simple random walk $\bar{X}_{i_{t}}$ are given by $\bar{P}_{i_{k,k\pm 1}}=\frac{1}{2}$ and $\bar{\mu}_{i}(k)=1/K$, respectively. Furthermore, the log-Soblev constant $\rho(\bar{P}_{i})$ is lower bounded by $\rho(\bar{P}_{i})\geq\frac{8\pi^{2}}{25K^{2}}$ (Diaconis and Saloff-Coste, 1996, Example 4.2). Thus, by definition the product chain $\bar{X}_{t}$ (Diaconis and Saloff-Coste, 1996, Sec. 2.5) has the following transition matrix:
and stationary distribution:
Furthermore, by Lemma $3.2$ in (Diaconis and Saloff-Coste, 1996) the log-Soblev constant $\rho(\bar{P})$ of the product chain $\{\bar{X}_{t}\}_{t\in\mathbb{N}}$ is lower bounded by:
$\rho(\bar{P})=\frac{1}{N}\min_{i\in\{1,\ldots,N\}}\rho(\bar{P}_{i})\geq\frac{8\pi^{2}}{25NK^{2}}.$ (19)
Next, note that there is a one-to-one mapping between the set $A^{N}$ and the set $\mathbb{Z}_{K}$ with $|\mathcal{A}|=A=K$ and thus a one-to-one mapping between the set $\mathcal{A}^{N}$ and the set $\mathbb{Z}_{K^{N}}$ with with $|\mathcal{A}^{N}|=A^{N}=K^{N}$. Therefore, we can assume that the Markov chains $\hat{X}_{t}$ and $\bar{X}_{t}$ operate on the same state space. By Equation (19) and with the following upper bounds on $M$ and $C$:
the log-Soblev constant $\rho(\hat{P})$ can be lower-bounded by:
$\rho(\hat{P})\geq\frac{1}{MC}\rho(\bar{P})\geq\frac{16\pi^{2}}{25NA^{3}},$
which concludes the proof.
Next, we proceed to prove Lemma 3.4.
Proof B.5.
of Lemma 3.4 The Markov chain $X_{t}$ with transition matrix $P$ defined in (7) is aperiodic and irreducible and thus a unique stationary distribution $\mu$ exists with $\mu^{t}=\mu^{0}P^{t}\rightarrow\mu$ for $t\rightarrow\infty$ from any initial distribution $\mu^{0}$. We define $\mu_{\max}$ and $\mu_{\min}$ as $\max_{a\in\mathcal{A}}\mu(a)\leq\mu_{\max}\leq 1$ and $\min_{a\in\mathcal{A}}\mu(a)\geq\mu_{\min}>0$, where $\mu_{\min}>0$ follows from the irreducibility of $X_{t}$.
Next, we consider the modified Markov chain $\{X^{*}_{t}\}_{t\in\mathbb{N}}$ with transition matrix $PP^{*}$ which is aperiodic and irreducible since $X_{t}$ is aperiodic and irreducible. Concretely, since $P$ contains self-loops, i.e., $P_{a,a}>0$, it follows that $PP^{*}$ contains self-loops, i.e.,
$PP^{*}_{a,a}=\sum_{a^{\prime}\in\mathcal{A}}P_{a,a^{\prime}}P^{*}_{a^{\prime},a}=\sum_{a^{\prime}\in\mathcal{A}}P_{a,a^{\prime}}\frac{\mu(a)P_{a,a^{\prime}}}{\mu(a^{\prime})}\geq P_{a,a}P_{a,a}>0,$
@@ Appendix B Convergence of log-linear learning
where we used that $P^{N}_{a,\tilde{a}}>0$ and $P_{a,a}>0$ for all $a,\tilde{a}\in\mathcal{A}^{N}$ as well as the identity $\mu(a)P^{*}_{a,\tilde{a}}=\mu(\tilde{a})P_{\tilde{a},a}$. It follows that $X^{*}_{t}$ is irreducible. Thus, for $X^{*}_{t}$ a unique stationary distribution $\mu^{*}$ exists. By Proposition $1.23$ in (Levin and Peres, 2017) the stationary distribution of $P^{*}$ is given by $\mu$ and thus the stationary distribution of $PP^{*}$ is given by $\mu$ since $\mu PP^{*}=\mu P^{*}=\mu$. Furthermore, the following holds for the transition matrix $PP^{*}$ :
@@
Next, we compute $M$ and $C$, defined bellow, of the Markov chains $X^{*}_{t}$ and $X_{t}$:
@@
From Corollary $2.15$ in (Diaconis and Saloff-Coste, 1996), it follows that the log-Soblev constant $\rho(PP^{*})$ is lower-bounded by:
$\rho(PP^{*})\geq\frac{1}{MC}\rho(P)\geq\frac{p_{\min}^{2}}{N}\rho(P).$ (20)
Next, we compare the Markov chain $X_{t}$ to the Markov chain $\hat{X}_{t}$ with transition matrix $\hat{P}$ specified in Equation (18) of Lemma B.3. To this end, we compute $M$ and $C$ of the Markov chains $X_{t}$ and $\hat{X}_{t}$:
@@
Thus, by Corollary $2.15$ in (Diaconis and Saloff-Coste, 1996) and by Lemma B.3 the log-Soblev constant $\rho(P)$ can be lower-bounded by:
$\rho(P)\geq\frac{1}{MC}\rho(\hat{P})\geq A^{N}A\mu_{\min}p_{\min}\rho(\hat{P})\geq\frac{16\pi^{2}A^{N}\mu_{\min}p_{\min}}{25NA^{2}}.$ (21)
Combining Equation (20) and (21), we conclude that the log-Soblev constant $\rho(PP^{*})$ is lower-bounded by:
@@
B.4 Modified log-linear learning in symmetric potential games
In the following, we consider a symmetric potential game and show that in this setting, we obtain convergence time guarantees to an $\epsilon$-efficient NE that depend polynomially on the number of players $N$ and on $1/\epsilon$. This improves the convergence time result provided in (Shah and Shin, 2010) which shows an exponential dependence on $1/\epsilon$.
A game is said to be symmetric if it satisfies the following definition.
Definition B.6.
A game is symmetric if for any permutation $\pi$ of $\{1,\ldots,N\}$ the following holds:
$U_{i}(a_{1},\ldots,a_{N})=U_{\pi(i)}(a_{\pi(1)},\ldots,a_{\pi(N)}).$
In other words, the utility of player $i$ depends only on how many players are playing each action $a\in\mathcal{A}$ and not on players’ identities. Thus, in a symmetric potential game, if $A<N$, the potential function $\Phi$ can be redefined in terms of a lower-dimensional function $\Phi:\Psi^{\mathcal{A}}\rightarrow[0,1]$, where
with cardinality $Y=|\Psi^{\mathcal{A}}|\leq(N+1)^{A-1}$. Then, for any $a=(a_{1},\ldots,a_{N})\in\mathcal{A}^{N}$ we have $\Phi(a)=\Phi(x(a))$ with $x(a)=(x_{1}(a),\ldots,x_{N}(a))$. Here $x_{j}(a)$ denotes the fraction of players that selected action $j\in A^{N}$, i.e., $x_{j}(a)=1/N|\{i\in[N]\>|\>a_{i}=j\}|$.
To obtain their results for symmetric potential games, Shah and Shin (2010) propose modified log-linear learning, a variant of log-linear learning. In modified log-linear learning every player $i$ has an independent exponential clock of rate $\alpha/z_{i}^{t}$, where $\alpha>0$ is a parameter and $z_{i}^{t}:=1/N|\{j\in[N]\>|\>a_{j}^{t}=a_{i}^{t}\}|$. [^8] The only change compared to log-linear learning is that in log-linear learning players have a fixed exponential clock rate of $1$. This means that the times between two consecutive clock-ticks are independent and distributed as the exponential distribution of mean $\alpha/z_{i}^{t}$. When the clock of player $i$ ticks, she is allowed to alter her current action. Player $i$ samples an action from her strategy $p_{i}^{t}\in\Delta(\mathcal{A})$ defined as in Equation (3). Modified log-linear learning induces an aperiodic and irreducible Markov chain on the lower-dimensional state space $\Psi^{\mathcal{A}}$ with stationary distribution $\mu_{m}\in\Delta(\Psi^{\mathcal{A}})$ given by (Shah and Shin, 2010, Lemma 2):
$\mu_{m}(x)=\frac{e^{\beta\Phi(x)}}{\sum_{\tilde{x}\in\Psi_{N}^{\mathcal{A}}}e^{\beta\Phi(\tilde{x})}}\quad\forall x\in\Psi^{\mathcal{A}}.$
In the following, we show that in a symmetric potential game, if all players adhere to modified log-linear learning an $\epsilon$-efficient NE is reached in time polynomial in $N$ and $1/\epsilon$.
Corollary B.7.

### v4

@@ 3.3 Convergence time in symmetric potential games
where $M=\max_{a\in\mathcal{A}^{N}}\frac{\mu(a)}{\hat{\mu}(a)}$ and $C=\max_{a\neq\tilde{a}:(P)_{a,\tilde{a}}\neq 0}\frac{\hat{\mu}(a)\hat{P}_{a,\tilde{a}}}{\mu(a)(P)_{a,\tilde{a}}}$.
Proof (Lemma 4.2)
Consider a Markov chain $X_{t}^{*}$ with transition matrix $PP^{*}$. Let $X_{t}$ be the Markov chain with transition matrix $P$ defined in Inequality (7). We will use Lemma 4.3 to obtain a lower bound on the log-Sobolev constant $\rho(PP^{*})$ in terms of $\rho(P)$.
Comparison of $X_{t}^{*}$ and $X_{t}$:
Note that $X_{t}$ is aperiodic and irreducible and thus a unique stationary distribution $\mu$ exists with $\mu^{t}=\mu^{0}P^{t}\rightarrow\mu$ for $t\rightarrow\infty$, where $\mu^{0}$ is any initial distribution. Furthermore, $\mu_{\min}>0$ follows from the irreducibility of $X_{t}$.
The Markov chain $X^{*}_{t}$ is also aperiodic and irreducible since $X_{t}$ is aperiodic and irreducible. Concretely, since $P$ contains self-loops, i.e., $P_{a,a}>0$, it follows that $PP^{*}$ contains self-loops:
$(PP^{*})_{a,a}$ $=\sum_{a^{\prime}\in\mathcal{A}}P_{a,a^{\prime}}P^{*}_{a^{\prime},a}=\sum_{a^{\prime}\in\mathcal{A}}P_{a,a^{\prime}}\frac{\mu(a)P_{a,a^{\prime}}}{\mu(a^{\prime})}$
$\geq P_{a,a}P_{a,a}>0,$
@@
where we used that $P^{N}_{a,\tilde{a}}>0$ and $P_{a,a}>0$ for all $a,\tilde{a}\in\mathcal{A}^{N}$ as well as the identity $\mu(a)P^{*}_{a,\tilde{a}}=\mu(\tilde{a})P_{\tilde{a},a}$. It follows that $X^{*}_{t}$ is irreducible. Thus, for $X^{*}_{t}$, a unique stationary distribution exists. More specifically $\mu$ is the stationary distribution of $PP^{*}$ since by [23, Proposition 1.23] the stationary distribution of $P^{*}$ is given by $\mu$ and since $\mu PP^{*}=\mu P^{*}=\mu$. Furthermore, the following holds for the transition matrix $PP^{*}$:
@@
Now, we apply Lemma 4.3 to lower-bound the log-Soblev constant $\rho(PP^{*})$ of $X_{t}^{*}$ in terms of the log-Sobolev constant $\rho(P)$ of $X_{t}$ as follows:
$\rho(PP^{*})\geq\frac{1}{MC}\rho(P)\geq\frac{p_{\min}^{2}}{N}\rho(P).$ (15)
where
@@
Next, we consider the Markov chain $\hat{X}_{t}$ with transition matrix $\hat{P}$ specified as $\hat{P}_{a,\tilde{a}}=\frac{1}{NA}\mathds{1}_{\tilde{a}\in\mathcal{N}(a)}$, where $\mathcal{N}(a)=\{\tilde{a}\in\mathcal{A}^{N}\>|\>\exists i\in[N]:\tilde{a}_{-i}=a_{-i}\}$.
Comparison of $X_{t}$ and $\hat{X}_{t}$:
Note that $\hat{X}_{t}$ is aperiodic and irreducible with stationary distribution $\hat{\mu}(a)=1/A^{N}$. This can be verified by checking the detailed balance equations given by $\hat{\mu}(a)\hat{P}_{a,\tilde{a}}=\hat{\mu}(\tilde{a})\hat{P}_{\tilde{a},a}$ for all $a,\tilde{a}\in A^{N}$.
[…]
@@
Thus, the log-Soblev constant $\rho(P)$ can be lower-bounded by:
[…]
Lastly, we consider the product chain $\bar{X}_{t}$ with $\bar{X}_{t}=\prod_{i=1}^{N}\bar{X}_{i_{t}}$ on the state space $\mathbb{Z}_{K^{N}}=\prod_{i=1}^{N}\mathbb{Z}_{K}$ with $\mathbb{Z}_{K}=\{1,\ldots,K\}$ and $K\geq 4$.
Comparison of $\hat{X}_{t}$ and $\bar{X}_{t}$:
Here, each $\{\bar{X}_{i,t}\}_{t\in\mathbb{N}}$ is a simple random walk on $\mathbb{Z}_{K}$ with transition matrix $\bar{P}_{i_{k,k\pm 1}}$ specified as:
$\bar{P}_{i,(k,k\pm 1)}=1/2\qquad\text{ for }2\leq k\leq K-1,$
$\bar{P}_{i,(K,1)}=\bar{P}_{i,(K,K-1)}=1/2,$
$\bar{P}_{i,(1,2)}=\bar{P}_{i,(1,K)}=1/2,$
and the stationary distribution $\bar{\mu}_{i}(k)$ of the simple random walk $\bar{X}_{i,t}$ is given by:
$\bar{\mu}_{i}(k)=\frac{1}{K},\quad\forall i\in\mathcal{N}.$
Thus, the product chain $\bar{X}_{t}$ has the following transition matrix [16, Sec. 2.5]:
and the stationary distribution:
Note that there is a one-to-one mapping between the set $\mathcal{A}$ and the set $\mathbb{Z}_{K}$ with $|\mathcal{A}|=A=K$ and thus a one-to-one mapping between the set $\mathcal{A}^{N}$ and the set $\mathbb{Z}_{K^{N}}$ with with $A^{N}=K^{N}$. Therefore, we can assume that the Markov chains $\hat{X}_{t}$ and $\bar{X}_{t}$ operate on the same state space. To this end, we compute $M$ and $C$ of the Markov chains $\hat{X}_{t}$ and $\bar{X}_{t}$:
[…]
$\rho(\hat{P})\geq\frac{1}{MC}\rho(\bar{P})\geq\frac{2}{A}\rho(\bar{P}).$
For the simple random walk $\bar{X}_{i,t}$ a bound on the log-Sobolev constant $\rho(\bar{P}_{i})$ is known with $\rho(\bar{P}_{i})\geq\frac{8\pi^{2}}{25K^{2}}$ [16, Example 4.2]. Then, by [16, Lemma 3.2], the log-Soblev constant $\rho(\bar{P})$ of the product chain $\bar{X}_{t}$ is lower bounded by:
$\rho(\bar{P})=\frac{1}{N}\min_{i\in\{1,\ldots,N\}}\rho(\bar{P}_{i})\geq\frac{8\pi^{2}}{25NK^{2}}.$
[…]
$\rho(\hat{P})\geq\frac{2}{A}\rho(\bar{P})\geq\frac{16\pi^{2}}{25NA^{3}}.$ (17)
Combination:
Combining Equations (15), (16), and (17), we conclude that the log-Sobolev constant $\rho(PP^{*})$ is lower-bounded by:
@@
This concludes the proof of Lemma 4.2.
5 Robustness of log-linear learning
In this section, we show a convergence time guarantee in settings where players have less feedback information, Subsection 5.1, and in settings where log-linear learning is subject to perturbations such as noisy utility observations, Subsection 5.2.
5.1 Reduced feedback
Log-linear learning requires players to observe their utilities for all possible actions given the other players’ actions. Having such full-information feedback when action sets are large can be demanding. Binary log-linear learning [1, 25] alleviates this limitation by requiring two-point feedback, reducing the needed feedback by a factor $A$ per round. We briefly review the binary log-linear learning rule.
$p_{i}^{t}(a_{i})=\begin{cases}\frac{e^{\beta U_{i}(a_{i},a_{-i}^{t-1})}}{e^{\beta U_{i}(a_{i}^{t-1},a_{-i}^{t-1})}+e^{\beta U_{i}(\tilde{a}_{i},a_{-i}^{t-1})}},&\text{for $a_{i}\in\{\tilde{a}_{i},a_{i}^{t-1}\}$.}\\ 0,&\text{otherwise.}\end{cases}$
Here, player $i$ can either repeat her action $a_{i}^{t-1}$ or play one other randomly sampled action $\tilde{a}_{i}$ rather than any action $a_{i}\in\mathcal{A}$ as in log-linear learning. Next, we derive the first finite-time convergence bound of binary log-linear learning to an $\epsilon$-efficient Nash equilibrium.
Theorem 5.1.